Derived the relation between the molarity and normality of x gram of suphuric acid dissolved in 100ml of water
1
Expert's answer
2020-08-06T03:58:04-0400

You can convert from molarity to normality using the following equation:

"Normality=Molarity\times\frac{Molecularmass}{Equivalent mass}"

For sulphuric acid "Normality=Molarity\times\frac{98}{49}=Molarity\times2"

Molarity of sulphuric acid = "\frac{chemical amount}{solutionvolume}=\frac{\frac{x}{98}}{\frac{x+100}{\rho}}"

where "\rho" - solution dencity

Normality of sulphuric acid = "=\frac{\frac{2x}{98}}{\frac{x+100}{\rho}}"